OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=5/6, (9,-4)

OpenStudy (anonymous):

y--4/x-9=5/6 y+4/x-9=5/6 6(y+4)=5(x-9) 6y+24=5x-45 6y=5x-69 y=5/6x-11.5

OpenStudy (anonymous):

11.5

OpenStudy (anonymous):

\[y-4=\frac{5}{6}(x-9)\]\[y-4=\frac{5}{6}x-\frac{5*9}{6}\]\[y=\frac{5}{6}x-\frac{45}{6}+4\]\[y=\frac{5}{6}x-\frac{45+(4*6)}{6}\]\[y=\frac{5}{6}x-\frac{69}{6}\]\[y=\frac{5}{6}x-\frac{23}{2}\]